A is 11 years old and B is 68 years old. in how many years will B be 4 times as old as A?
Found 3 solutions by Alan3354, Theo, MathTherapy:
Answer by Alan3354(69443)

A is 11 years old and B is 68 years old. in how many years will B just 4 times as old as B?
----------
The difference is 57 years, and the difference is 3 times A's age.

Answer by Theo(13342)

A is 11 years old.
B is 68 years old.

let the number of years be equal to x.

in x years, B will be 68 + x years old.
in x years, A will be 11 + x years old.
in x years, B will be 4 * as old as A.

your equation is:
B + x = 4 * (A + x)
simplify to get:
B + x = 4 * A + 4 * x
subtract 4 * x from both sides of the equation to get:
B + x - 4 * x = 4 * A
combine like terms to get:
B - 3 * x = 4 * A
subtract 4 * A from both sides of the equation and add 3 * x to both sides of the equation to get:
B - 4 * A = 3 * x
since B = 68 and A = 11, this equation becomes:
68 - 4 * 11 = 3 * x
simplify to get:
68 - 44 = 3 * x
combine like terms to get:
24 = 3 * x
solve for x to get:
x = 24/3 = 8

your solution is B will be 4 * as old as A in 8 years.
in 8 years, B will be 68 + 8 = 76
in 8 years, A will be 11 + 8 = 19
4 * 19 = 76, confirming the solution is correct.


Answer by MathTherapy(10552)


A is 11 years old and B is 68 years old. in how many years will B be 4 times as old as A?
With "y" being the number of years, we get the following SIMPLE equation: 68 + y = 4(11 + y)
Solve this to get y = 8, so in 8 YEARS.
That's all!!